# Scaffold-BPE: Enhancing Byte Pair Encoding for Large Language Models with Simple and Effective Scaffold Token Removal

### v2

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have become a burgeoning paradigm in handling a broad array of Natural Language Processing (NLP) tasks. The tokenization process in most modern LLMs (Radford et al. 2019; Brown et al. 2020; Rae et al. 2021; Zhang et al. 2022; Biderman et al. 2023; Touvron et al. 2023a; Yang et al. 2023; Achiam et al. 2023; Dubey et al. 2024) employs Byte Pair Encoding (BPE) (Sennrich et al. 2015), a method that was originally designed for data compression (Gage 1994). BPE consists of two main stages. In the training stage, BPE iteratively merges the most frequent pairs of bytes or characters in a dataset into a new token and adds it to the vocabulary until a desired vocabulary size is reached. And in the encoding stage, the vocabulary is utilized to represent any text. The adoption of BPE in LLMs is driven by its capability to decompose words into smaller, manageable subword units, thus avoiding out-of-vocabulary words, facilitating flexible and semantically complete representations of input data. Actually, BPE has also been widely used in traditional NLP tasks, like machine translation (Provilkov et al. 2019; Xu et al. 2020), sentence classification (Liu et al. 2019; He et al. 2020b) and summarization (Wu et al. 2021; Xu et al. 2022).
@@
Since its inception, BPE has undergone various modifications to better suit the needs of complex NLP tasks, including identifying the optimal vocabulary size for various tasks (Xu et al. 2020; Gutierrez-Vasques et al. 2021), optimizing the encoding paths of tokens to achieve subword regularization (Provilkov et al. 2019; He et al. 2020a), etc.
@@
However, existing studies have overlooked an inherent limitation in the BPE method: the iterative merging process can lead to an imbalance in token frequencies by including low-frequency tokens in vocabulary. For example, as illustrated in Figure 1, in the commonly used Pile dataset (Gao et al. 2020) that is tokenized by the original BPE method of 32K vocabulary size (as LLaMA series (Touvron et al. 2023a, b)), the token “zona” mostly appears as a component of the token “Arizona” rather than as an independent, high-frequency token. Despite its lower standalone frequency, BPE includes “zona” in the final vocabulary because it is the “intermediate token” to derive the frequent token “Arizona”. We define such intermediate tokens that are crucial for constructing longer frequent tokens but do not appear frequently on their own as Scaffold Tokens. Note that not all subwords are simply scaffold tokens. For example, “ing” is not identified as a scaffold token, as there are many words containing “ing” but are not tokens in the vocabulary. For example, “connecting” is represented as “connect”+“ing” (2 tokens). Such words help to keep “ing” a frequent token. Therefore, “ing” is not a scaffold token. According to our proposed Scaffold-BPE method, the 32K vocabulary contains about 6.07% of scaffold tokens.
As depicted in Figure 2, a natural frequency imbalance arises between these scaffold tokens and actual high-frequency tokens. Prior studies (Lin et al. 2017; Su et al. 2023) have highlighted that such disparities in token frequencies can result in imbalanced learning difficulties across different tokens. Scaffold tokens, due to their lower individual occurrence frequencies, are notably harder to learn for models.
@@
We conduct extensive experiments on language modeling tasks. Results on 9 widely used language modeling benchmarks demonstrate that Scaffold-BPE consistently outperforms the original BPE. Besides, even when extended to machine translation tasks, Scaffold-BPE proves highly effective. Furthermore, we show that Scaffold-BPE is orthogonal to existing modifications on BPE, like BPE-Dropout (Provilkov et al. 2019) and can be combined with them to achieve further improvements.
@@ 2.1 Language Models
Language models are designed to maximize the likelihood of a token sequence. Following GPT-3 (Brown et al. 2020), which features 175 billion parameters and demonstrates versatility across a wide range of applications, there has been a significant push towards developing large generative language models like Gopher (Rae et al. 2021), PaLM (Chowdhery et al. 2023), GaLM (Du et al. 2022), OPT (Zhang et al. 2022), and LLaMA (Touvron et al. 2023a). Such a surge in development has greatly advanced the fields of natural language understanding and generation.
@@ 2.2 Byte Pair Encoding
Early neural models had difficulty managing rare words due to limited vocabulary sizes. BPE (Sennrich et al. 2015) effectively addresses that by generating a subword vocabulary. Initially, a corpus is split into characters or bytes, which act as initial tokens. The algorithm iteratively finds the most frequent token pair in the sequence, merges them into a new token, and adds it to the vocabulary until it reaches a predetermined size. The vocabulary is then utilized during the encoding phase to represent any text. Recent advancements like BPE-dropout (Provilkov et al. 2019) and optimal vocabulary size search (Xu et al. 2020; Gowda et al. 2020; Salesky et al. 2020) continue to enrich BPE developments.
@@ 3 Methodology
0:  Text Corpus $C$, Desired Vocabulary Size $N$
@@
2:  Split $C$ into a list of characters/bytes (denoted as $L$)
@@
8:        // $t$ may be a previously marked scaffold token
@@
15:     /****** Scaffold-BPE Modification Begins ******/
@@
23:     /****** Scaffold-BPE Modification Ends ******/
@@
3.1 Training Process
$(a,b)=\arg\max_{(x,y)\in L}f((x,y))$ (1)
The iterative process of identifying the most frequent token pair $(a,b)$ can be accelerated using a priority queue $Q$. At the beginning of the training process, all token pairs in $L$ are pushed into $Q$ with a descending order of frequency. And after the token pair $(a,b)$ is merged into $t$ in each iteration, BPE updates the frequencies and the ranks of token pairs related to all indexed occurrences of $(a,b)$. For instance, given $(a,b)$ in a context of “$\ldots,u,a,b,v,\ldots$” in $L$, when $(a,b)$ is replaced with $t$, the frequency of $(u,a)$ or $(b,v)$ would decrease by 1, and meanwhile that of $(u,t)$ or $(t,v)$ would increase by 1. With the occurrences of all token pairs being indexed, there is no need to scan $L$ again and re-count the frequencies of all candidate token pairs for a new iteration. After updating the adjacent token pairs related to $(a,b)$ (i.e, $t$), the frequencies of token pairs like $(u,a)$ or $(b,v)$ would be updated in $Q$, and meanwhile the new candidate token pairs $(u,t)$ and $(t,v)$ would also be pushed into $Q$ with their frequencies.
$V=\{t\in E\mid\text{Scaffold}(t)=False\}$ (2)
$S=\{t\in E\mid\text{Scaffold}(t)=True\}$ (3)
@@ 3.1 Training Process
3.2 Encoding Process
$t=\arg\min_{t\in E}rank_{E}(t)$ (5)
Figure 4: The encoding process (i.e., Scaffolding and Demolishing) for the word “zona” and “Arizona”.
4 Experiments
4.1 Experimental Setup
Datasets.
Our models are trained on the Pile (Gao et al. 2020) dataset, an 825.18 GiB English text dataset designed for training LLMs. The data distribution for our model training is identical to that of the original work (Gao et al. 2020).
Tokenizer.
We train two 32K vocabularies (size applied by LLaMA series (Touvron et al. 2023a, b)) using the original BPE and Scaffold-BPE, respectively. Similar to GPT-2 (Radford et al. 2019), pre-tokenization was employed to prevent the merging of tokens from different character categories. And following (Touvron et al. 2023a), we split numbers into individual digits.
Model.
We train three language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures of the 468M and the 1.2B models are identical to those of the 410M and the 1.0B models outlined in Pythia (Biderman et al. 2023). The minor differences in parameter sizes are attributed to the variations in vocabulary size in the embedding and output layer. As for the 6.7B model, its architecture is identical to LLaMA-7B (Touvron et al. 2023a).
Training.
Following the pretraining settings of previous works (Xie et al. 2023; Su et al. 2024; Xiong et al. 2024) and limited by our computation budget, by default all models are pretrained with 100B tokens. Note that the volume of corresponding text data contained in an equal amount of tokens is slightly different between the two tokenizers. Considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, we compare experiments in the setting of an equal amount of training tokens. In the Section 4.3, we further analyze both tokenizers in the setting of an equal amount of training text volume.
4.2 Experimental Results
Common Sense Reasoning.
Our analysis incorporates 7 benchmarks recognized for evaluating common sense reasoning, including BoolQ (Clark et al. 2019), HellaSwag (Zellers et al. 2019), OpenBookQA (Mihaylov et al. 2018), PIQA (Bisk et al. 2020), SIQA (Sap et al. 2019), StoryCloze (Mostafazadeh et al. 2016), and Winogrande (Sakaguchi et al. 2021). We present the performance of all models in terms of average accuracy in 0-shot and 5-shot settings.
@@ 4.2 Experimental Results
Closed Book Question Answering.
For the task of closed book question answering (Brown et al. 2020; Touvron et al. 2023a), we evaluate the performance of the largest 6.7B-parameter models with different tokenizers on 2 benchmark datasets, i.e., TriviaQA (Joshi et al. 2017) and WebQuestions (Berant et al. 2013). We report the exact match performance for the zero-shot and few-shot settings in Table 2. It can be seen that the model trained with the proposed Scaffold-BPE can achieve a 3.23pp improvement on TriviaQA and a 1.33pp improvement on WebQuestions, with both $p$-values less than 0.01. All results above demonstrate that Scaffold-BPE can enhance model performance across different types of downstream tasks.
@@ 4.3 Discussion
Various Vocabulary Size.
Depending on the size of the training corpus, the diversity of the languages, the size of the model, and the types of tasks, different vocabulary sizes are set in practice. Therefore, to validate the robustness of Scaffold-BPE across various vocabulary sizes, in addition to the 32K vocabulary (Touvron et al. 2023a), we also trained two vocabularies sized at 64K (Baichuan 2023b, a) and 128K (Yang et al. 2023). The experimental setup is identical to that of the 468M-parameter model mentioned before.
@@
More Training Tokens.
According to the Scaling Law, the loss scales as a power-law with model size, dataset size, and the amount of training computation (Kaplan et al. 2020). To demonstrate the effectiveness of our Scaffold-BPE with more training tokens, we continue training the 468M models up to 300B tokens (Zhang et al. 2022; Biderman et al. 2023).
@@
Applicable for Other Tasks, Languages, Model Architectures and Compatible with Other BPE Enhancements.
Although the development of LLMs is burgeoning, some applications still prefer using conventional models due to their lower training and inference costs. In the NLP field, BPE was initially combined with transformer models and applied to machine translation tasks (Sennrich et al. 2015), which typically face an open vocabulary challenge and involve substantial textual variations between two languages. Therefore, to validate the versatility of the Scaffold-BPE method, we additionally conduct evaluations on machine translation tasks with identical experimental setup on WMT’14 En-De and En-Fr dataset in the prior work (Ott et al. 2018).
@@
Moreover, previous experiments on language modeling tasks are carried out on the decoder-only architecture. For the machine translation tasks, we utilize the encoder-decoder architecture (Vaswani et al. 2017). The exceptional performance of Scaffold-BPE confirms its architecture insensitivity, indicating its applicability across a wider range of neural network architectures.
Finally, Scaffold-BPE is orthogonal to and can be combined with existing enhancements to BPE, like BPE-Dropout (Provilkov et al. 2019). As shown in Table 5, Scaffold-BPE with BPE-Dropout achieves further improvements on BLEU, well indicating the compatibility of Scaffold-BPE.
@@
Higher Compression Rate.
Experiments under Same Corpus Size
@@ A.1 Higher Entropy, Lower Redundancy
Scaffold-BPE can alleviate the imbalance in token frequency, which can lead to an increase in information entropy. We measure Shannon Entropy and Redundancy (Gutierrez-Vasques et al. 2021) over token representations of texts obtained with the original BPE and our Scaffold-BPE. Both take as input a text $T$ with a vocabulary of (normal) tokens $V=\{t_{1},t_{2},...,t_{V}\}$ of size $|V|$.
(a) The Original BPE
(b) The Scaffold-BPE
@@
As shown in Table 8, taking the 32K vocabulary as an example, our Scaffold-BPE can encode Pile dataset (Gao et al. 2020) with higher Entropy and lower Redundancy. Consequently, tokens in the vocabulary of our Scaffold-BPE have more balanced appearing probabilities. According to Su et al. (2023), our vocabulary with balanced token occurrences mitigates the learning imbalance problem, resulting in more sufficient learning towards the text corpus, thus achieving better performance.

### v3

@@ Abstract
Scaffold-BPE: Enhancing Byte Pair Encoding for Large Language Models with Simple and Effective Scaffold Token Removal
Haoran Lian1, Yizhe Xiong2,3, Jianwei Niu1, Shasha Mo1, Zhenpeng Su4, Zijia Lin2, Hui Chen2,3, Peng Liu5, Jungong Han2, Guiguang Ding2,3, 1Beihang University, 2Tsinghua University, 3BNRist, 4Chinese Academy of Sciences, 5Beijing Institute Of Technology Correspondence: niujianwei@buaa.edu.cn
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have become a burgeoning paradigm in handling a broad array of Natural Language Processing (NLP) tasks. The tokenization process in most modern LLMs Radford et al. (2019); Brown et al. (2020); Rae et al. (2021); Zhang et al. (2022); Biderman et al. (2023); Touvron et al. (2023a); Yang et al. (2023); Achiam et al. (2023); Dubey et al. (2024) employs Byte Pair Encoding (BPE) Sennrich et al. (2015), a method that was originally designed for data compression (Gage, 1994). BPE consists of two main stages. In the training stage, BPE iteratively merges the most frequent pairs of bytes or characters in a dataset into a new token and adds it to the vocabulary until a desired vocabulary size is reached. And in the encoding stage, the vocabulary is utilized to represent any text. The adoption of BPE in LLMs is driven by its capability to decompose words into smaller, manageable subword units, thus avoiding out-of-vocabulary words, facilitating flexible and semantically complete representations of input data. Actually, BPE has also been widely used in traditional NLP tasks, like machine translation (Provilkov et al., 2019; Xu et al., 2020), sentence classification (Liu et al., 2019; He et al., 2020b) and summarization (Wu et al., 2021; Xu et al., 2022).
@@
Since its inception, BPE has undergone various modifications to better suit the needs of complex NLP tasks, including identifying the optimal vocabulary size for various tasks Xu et al. (2020); Gutierrez-Vasques et al. (2021), optimizing the encoding paths of tokens to achieve subword regularization Provilkov et al. (2019); He et al. (2020a), etc.
@@
However, existing studies have overlooked an inherent limitation in the BPE method: the iterative merging process can lead to an imbalance in token frequencies by including low-frequency tokens in vocabulary. For example, as illustrated in Figure 1, in the commonly used Pile dataset (Gao et al., 2020) that is tokenized by the original BPE method of 32K vocabulary size (as LLaMA series Touvron et al. (2023a, b)), the token “zona” mostly appears as a component of the token “Arizona” rather than as an independent, high-frequency token. Despite its lower standalone frequency, BPE includes “zona” in the final vocabulary because it is the “intermediate token” to derive the frequent token “Arizona”. We define such intermediate tokens that are crucial for constructing longer frequent tokens but do not appear frequently on their own as Scaffold Tokens. Note that not all subwords are simply scaffold tokens. For example, “ing" is not identified as a scaffold token, as there are many words containing “ing" but are not tokens in the vocabulary. For example, “connecting" is represented as “connect"+“ing" (2 tokens). Such words help to keep “ing” a frequent token. Therefore, “ing" is not a scaffold token. According to our proposed Scaffold-BPE method, the 32K vocabulary contains about 6.07% of scaffold tokens.
As depicted in Figure 2, a natural frequency imbalance arises between these scaffold tokens and actual high-frequency tokens. Prior studies Lin et al. (2017); Su et al. (2023) have highlighted that such disparities in token frequencies can result in imbalanced learning difficulties across different tokens. Scaffold tokens, due to their lower individual occurrence frequencies, are notably harder to learn for models.
@@
We conduct extensive experiments on language modeling tasks. Results on 9 widely used language modeling benchmarks demonstrate that Scaffold-BPE consistently outperforms the original BPE. Besides, even when extended to machine translation tasks, Scaffold-BPE proves highly effective. Furthermore, we show that Scaffold-BPE is orthogonal to existing modifications on BPE, like BPE-Dropout Provilkov et al. (2019) and can be combined with them to achieve further improvements.
@@ 2.1 Language Models
Language models are designed to maximize the likelihood of a token sequence. Following GPT-3 (Brown et al., 2020), which features 175 billion parameters and demonstrates versatility across a wide range of applications, there has been a significant push towards developing large generative language models like Gopher (Rae et al., 2021), PaLM (Chowdhery et al., 2023), GaLM (Du et al., 2022), OPT (Zhang et al., 2022), and LLaMA (Touvron et al., 2023a). Such a surge in development has greatly advanced the fields of natural language understanding and generation.
@@ 2.2 Byte Pair Encoding
Early neural models had difficulty managing rare words due to limited vocabulary sizes. BPE Sennrich et al. (2015) effectively addresses that by generating a subword vocabulary. Initially, a corpus is split into characters or bytes, which act as initial tokens. The algorithm iteratively finds the most frequent token pair in the sequence, merges them into a new token, and adds it to the vocabulary until it reaches a predetermined size. The vocabulary is then utilized during the encoding phase to represent any text. Recent advancements like BPE-dropout (Provilkov et al., 2019) and optimal vocabulary size search (Xu et al., 2020; Gowda et al., 2020; Salesky et al., 2020) continue to enrich BPE developments.
@@ 3 Methodology
3.1 Training Process
$(a,b)=\arg\max_{(x,y)\in L}f((x,y))$ (1)
The iterative process of identifying the most frequent token pair $(a,b)$ can be accelerated using a priority queue $Q$. At the beginning of the training process, all token pairs in $L$ are pushed into $Q$ with a descending order of frequency. And after the token pair $(a,b)$ is merged into $t$ in each iteration, BPE updates the frequencies and the ranks of token pairs related to all indexed occurrences of $(a,b)$. For instance, given $(a,b)$ in a context of “$\ldots,u,a,b,v,\ldots$" in $L$, when $(a,b)$ is replaced with $t$, the frequency of $(u,a)$ or $(b,v)$ would decrease by 1, and meanwhile that of $(u,t)$ or $(t,v)$ would increase by 1. With the occurrences of all token pairs being indexed, there is no need to scan $L$ again and re-count the frequencies of all candidate token pairs for a new iteration. After updating the adjacent token pairs related to $(a,b)$ (i.e, $t$), the frequencies of token pairs like $(u,a)$ or $(b,v)$ would be updated in $Q$, and meanwhile the new candidate token pairs $(u,t)$ and $(t,v)$ would also be pushed into $Q$ with their frequencies.
@@ 3.1 Training Process
0:  Text Corpus $C$, Vocabulary Size $N$
@@
2:  Split $C$ into a characters/bytes list $L$
@@
8:        /* $t$ may be a previously marked scaffold token */
@@
15:     /** Scaffold-BPE Modification Begins **/
@@
23:     /** Scaffold-BPE Modification Ends **/
@@
$V=\{t\in E\mid\text{Scaffold}(t)=False\}$ (2)
$S=\{t\in E\mid\text{Scaffold}(t)=True\}$ (3)
@@
3.2 Encoding Process
$t=\arg\min_{t\in E}rank_{E}(t)$ (5)
Figure 4: The encoding process (i.e., Scaffolding and Demolishing) for the word “zona" and “Arizona".
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets.
Our models are trained on the Pile (Gao et al., 2020) dataset, an 825.18 GiB English text dataset designed for training LLMs. The data distribution for our model training is identical to that of the original work Gao et al. (2020).
4.1.2 Tokenizer.
We train two 32K vocabularies (size applied by LLaMA series Touvron et al. (2023a, b)) using the original BPE and Scaffold-BPE, respectively. Similar to GPT-2 (Radford et al., 2019), pre-tokenization was employed to prevent the merging of tokens from different character categories. And following Touvron et al. (2023a), we split numbers into individual digits.
4.1.3 Model.
We train three language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures of the 468M and the 1.2B models are identical to those of the 410M and the 1.0B models outlined in Pythia Biderman et al. (2023). The minor differences in parameter sizes are attributed to the variations in vocabulary size in the embedding and output layer. As for the 6.7B model, its architecture is identical to LLaMA-7B Touvron et al. (2023a).
4.1.4 Training.
Following the pretraining settings of previous works Xie et al. (2023); Su et al. (2024); Xiong et al. (2024) and limited by our computation budget, by default all models are pretrained with 100B tokens. Note that the volume of corresponding text data contained in an equal amount of tokens is slightly different between the two tokenizers. Considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, we compare experiments in the setting of an equal amount of training tokens. In the Section 4.3.5, we further analyze both tokenizers in the setting of an equal amount of training text volume.
4.2 Experimental Results
@@
4.2.1 Common Sense Reasoning.
Our analysis incorporates 7 benchmarks recognized for evaluating common sense reasoning, including BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), StoryCloze (Mostafazadeh et al., 2016), and Winogrande (Sakaguchi et al., 2021). We present the performance of all models in terms of average accuracy in 0-shot and 5-shot settings.
@@
4.2.2 Closed Book Question Answering.
For the task of closed book question answering Brown et al. (2020); Touvron et al. (2023a), we evaluate the performance of the largest 6.7B-parameter models with different tokenizers on 2 benchmark datasets, i.e., TriviaQA (Joshi et al., 2017) and WebQuestions (Berant et al., 2013). We report the exact match performance for the zero-shot and few-shot settings in Table 2. It can be seen that the model trained with the proposed Scaffold-BPE can achieve a 3.23pp improvement on TriviaQA and a 1.33pp improvement on WebQuestions, with both $p$-values less than 0.01. All results above demonstrate that Scaffold-BPE can enhance model performance across different types of downstream tasks.
@@ 4.3 Discussion
4.3.1 Various Vocabulary Size.
Depending on the size of the training corpus, the diversity of the languages, the size of the model, and the types of tasks, different vocabulary sizes are set in practice. Therefore, to validate the robustness of Scaffold-BPE across various vocabulary sizes, in addition to the 32K vocabulary (Touvron et al., 2023a), we also trained two vocabularies sized at 64K (Baichuan, 2023b, a) and 128K (Yang et al., 2023). The experimental setup is identical to that of the 468M-parameter model mentioned before.
@@
4.3.2 More Training Tokens.
According to the Scaling Law, the loss scales as a power-law with model size, dataset size, and the amount of training computation Kaplan et al. (2020). To demonstrate the effectiveness of our Scaffold-BPE with more training tokens, we continue training the 468M models up to 300B tokens Zhang et al. (2022); Biderman et al. (2023).
@@
4.3.3 Applicable for Other Tasks, Languages, Model Architectures and Compatible with Other BPE Enhancements.
Although the development of LLMs is burgeoning, some applications still prefer using conventional models due to their lower training and inference costs. In the NLP field, BPE was initially combined with transformer models and applied to machine translation tasks Sennrich et al. (2015), which typically face an open vocabulary challenge and involve substantial textual variations between two languages. Therefore, to validate the versatility of the Scaffold-BPE method, we additionally conduct evaluations on machine translation tasks with identical experimental setup on WMT’14 En-De and En-Fr dataset in the prior work Ott et al. (2018).
@@
Moreover, previous experiments on language modeling tasks are carried out on the decoder-only architecture. For the machine translation tasks, we utilize the encoder-decoder architecture Vaswani et al. (2017). The exceptional performance of Scaffold-BPE confirms its architecture insensitivity, indicating its applicability across a wider range of neural network architectures.
Finally, Scaffold-BPE is orthogonal to and can be combined with existing enhancements to BPE, like BPE-Dropout Provilkov et al. (2019). As shown in Table 5, Scaffold-BPE with BPE-Dropout achieves further improvements on BLEU, well indicating the compatibility of Scaffold-BPE.
@@
4.3.4 Higher Compression Rate.
4.3.5 Experiments under Same Corpus Size
@@ A.1 Higher Entropy, Lower Redundancy
Scaffold-BPE can alleviate the imbalance in token frequency, which can lead to an increase in information entropy. We measure Shannon Entropy and Redundancy Gutierrez-Vasques et al. (2021) over token representations of texts obtained with the original BPE and our Scaffold-BPE. Both take as input a text $T$ with a vocabulary of (normal) tokens $V=\{t_{1},t_{2},...,t_{V}\}$ of size $|V|$.
@@
As shown in Table 8, taking the 32K vocabulary as an example, our Scaffold-BPE can encode Pile dataset Gao et al. (2020) with higher Entropy and lower Redundancy. Consequently, tokens in the vocabulary of our Scaffold-BPE have more balanced appearing probabilities. According to Su et al. (2023), our vocabulary with balanced token occurrences mitigates the learning imbalance problem, resulting in more sufficient learning towards the text corpus, thus achieving better performance.